# Canonical Electrodynamics in Ashtekar-Barbero variables

### v1

@@ 3.1 Weyl Fermions
3.2 Dirac fermions
Before analyzing the QED framework (i.e., the case involving both fermions and photons and their interaction), it is necessary to study Dirac fermions. All electrically charged fermions are described by Dirac spinors. In contrast, Majorana fermions —being their own antiparticles— must be electrically neutral. Since QED describes interactions between charged particles mediated by photon exchange, the theory inherently involves Dirac fermions. Consequently, we adopt the four-component representation of fermionic fields Pal:2011pb ; Schwartz:2014sze . For further details, refer to Appendix A.4.
The starting point is the following Dirac action
$S_{F}=\int d^{4}x|e|\Bigl{\{}\frac{i}{2}\left(e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\right)-m\overline{\Psi}\Psi\Bigr{\}}\,.$ (66)
where the covariant derivatives are defined in Eqs. 181-182. Since the mass term does not contain any derivatives, performing the same computation as in the massless case reveals that the only modification appears in the Hamiltonian constraint, as we will see below. Thus, the constraints for a system in which vacuum is minimally coupled to a massive fermionic field are
$\mathcal{G}_{i}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\,,$ (67)
$\mathcal{S}_{i}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}\,J_{i}\,,$ (68)
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\Bigr{\}}$ (69)
$\mathcal{H}_{a}=$ $\frac{1}{\kappa\beta}\left(E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\right)-\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\mathfrak{D}_{a}\Psi+\overline{\mathfrak{D}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (71)
In the previous expressions we have once again suppressed the superscript “G+F”.
Following the same steps as in the masseless case, we find that the compatibility condition remains the same as in the Weyl fermions case. As a result, we obtain the same modified spin connection which leads to the same new Ashtekar connection as given in Eq.(41). Hence, writing the fermionic convariant derivatives in terms of the new Ashtekar connection as in Eqs. (47)-(49) leads to the following constraints
$\mathcal{H}$ $=\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\right)-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+\sqrt{\text{det}\,h}\,m\,\overline{\Psi}\Psi\,,$ (73)
$\mathcal{H}_{a}$ $=\frac{1}{\beta\kappa}E_{j}^{b}\left(\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon_{kl}^{j}K_{a}^{k}K_{b}^{l}\right)+\frac{i}{2}\left(\overline{\Psi}\gamma^{0}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\gamma^{0}\Psi\right)$
$-\frac{\beta}{2}\sqrt{\text{det}\,h}K_{a}^{i}J_{i}\,,$ (74)
where, as mentioned before, the only difference with respect to the masseless case is the presence of a mass term in the Hamiltonian constraint.
We can once again perform a consistency check by comparing the EOM derived from the Euler-Lagrange equations with the ones derived from the Hamiltonian in the above form. The Euler-Lagrange equations yield gives
$ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma_{0}\Psi-m\Psi=0.$ (75)
By following the same steps as in the massless case, it is straightforward to show that the expression for the fermionic EOM in the Hamiltonian formulation matches this expression.
In contrast to the massless case, it is not convenient to use the bispinor notation and express the constraints in terms of the Pauli basis (see Appendix A.4 for further details on QED and chirality). We will instead present the constraints in terms of the half-density fields as follows
$+\frac{i}{2}\beta e_{i}^{a}D_{a}^{(A)}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{8}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}(1+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)-im\pi\gamma^{0}\,,$ (78)
4 QED
@@ 4.1 The full Hamiltonian
In this section, we examine a system in which photons —denoted by a $\gamma$ with no indices— are coupled to a fermionic field, and both fields are minimally coupled to gravity. Unlike the standard formulation of QED where the background spacetime is fixed, here the background is allowed to be dynamical. To maintain clarity and avoid an overload of calculations at once, we begin by analyzing the QED action in curved spacetime using modified Ashtekar-Barbero variables, an approach which, to the best of our knowledge, has not been previously explored in the literature. After deriving the QED contributions to the constraints, we will then proceed to minimally couple the full system to the gravitational field.
The QED action in curved spacetime takes the form:
$S_{\text{QED}}=$ $S_{F}+S_{\gamma}$ (80)
@@
×$\times(\nabla_{c}A_{d}-\nabla_{d}A_{c})$
@@
×$\times\left(\dot{A}_{d}-\nabla_{d}(t^{c}A_{c})-N^{c}F_{cd}\right)\,.$
@@
Let us now perform the same analysis for the full QED action $S_{\text{QED}}$ in (80). Beginning with the contributions involving the time vector $t^{a}$, we identify the following terms
@@
We observe that the first integral corresponds to the symplectic term, while the second integral contains two Gauss constraints, each multiplied by a Lagrange multiplier, respectively $\omega_{t}^{ij}$ and $A_{t}$. One of these constraints depends solely on internal indices and represents the fermionic contribution to the vacuum Gauss constraint. The other involves only spacetime indices —specifically, the spatial ones— and reduces to Gauss’ Law in the absence of the fermionic field. Hence, the full QED contribution to the Gauss constraint is
$\Lambda\mathcal{G}^{\text{QED}}=\omega_{t}{}^{ij}\,\frac{i}{2}\sqrt{\text{det}\,h}\,\Psi^{\dagger}\sigma_{ij}\Psi+A_{t}\left(\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\Psi^{\dagger}\Psi\right)\,.$ (91)
@@
Next, considering the terms involving the lapse function $N$, we obtain the QED Hamiltonian constraint as
$\begin{split}\mathcal{H}^{\text{QED}}&=\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{i}\Psi\right)\\ &+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,.\end{split}$ (92)
Finally, the terms multiplied by the shift vector $N^{a}$ yield the QED diffeomorphism constraint
$\mathcal{H}_{a}^{\text{QED}}=F_{ab}\pi^{b}+\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (93)
@@
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{QED}}=$ $\left(\beta\omega_{t}{}^{0i}-\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{t}{}^{jk}\right)\Bigl{[}\frac{1}{\kappa\beta}\left(\partial_{a}E_{i}^{a}+\epsilon_{ip}{}^{q}A_{a}^{p}E_{q}^{a}\right)+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\Bigr{]}$
@@
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
@@
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
@@
$-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+qE_{i}^{a}A_{a}\mathcal{J}^{i}+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,,$ (101)
@@
$\begin{split}ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi&-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi\\ &-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\Psi-m\Psi-qA_{a}e_{0}^{a}\gamma^{0}\Psi-qA_{a}e_{i}^{a}\gamma^{i}\Psi=0\,,\end{split}$ (103)
which is consistent with the well-known QED EOM for fermions. On the other hand, the EOM for the photon field can be obtained by extremizing the full action $S_{\text{Full}}$ with respect to the electromagnetic 4-vector potential $A_{a}$. This yields a background-independent form of Maxwell’s equations. More concretely, the photonic Euler-Lagrange equation reads
$\frac{\delta\mathcal{L}^{\text{QED}}}{\delta A_{b}}-\nabla_{c}\frac{\delta\mathcal{L}^{\text{QED}}}{\delta\left(\nabla_{c}A_{b}\right)}=0\,,$ (104)
@@
which is consistent with standard QED theory.
@@
$(t^{a}-N^{a})$ $\Bigl{[}i\gamma^{0}D_{a}^{(A)}\xi-\frac{\beta}{2}K_{a}^{i}\gamma_{5}\gamma_{i}\xi-qA_{a}\gamma^{0}\xi\Bigr{]}-N\Bigl{[}ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi$
@@
Finally, once again we impose the compatibility condition and observe that the $\mathcal{S}_{i}$ constraint is proportional to the Gauss constraint arising from the fermionic modification of the vacuum. As a result, $\mathcal{S}_{i}$ can be discarded. Therefore, the remaining (first class) constraints are the Gauss constraint derived in the previous section and the one associated with the QED sector, together with the Hamiltonian and the diffeomorphism constraints. Moreover, since we are working with Dirac spinors —and more generally within the framework of QED— the considerations regarding chirality remain applicable. Consequently, the final full constraints, expressed in terms of the modified Ashtekar-Barbero connection and the half-density fermionic field, are
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{QED}}=$ $\Lambda^{i}\Bigl{[}\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{i}{2}\frac{\beta^{2}}{1+\beta^{2}}\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi\Bigr{]}+A_{t}\left(\nabla_{a}\pi^{a}+iq\pi\xi\right)\,,$ (110)
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}\right)$
@@
$-\frac{1}{8}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}(1+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-im\pi\gamma^{0}$
@@ 4.2 Parity transformation
In this final subsection, we analyze the behavior of the full system under parity transformation. A similar study was conducted in Das:2008r for a system where gravity is coupled solely to a fermionic field. Here, we extend the analysis to the full system, where the gravitational field is minimally coupled to both fermionic and photonic fields, providing clearer, more detailed, and more readable final results.
@@
$+N\left[\frac{1}{\beta\sqrt{\text{det}\,h}}\left(E_{p}^{c}E_{m}^{a}A_{a}^{m}-E_{m}^{c}E_{p}^{a}A_{a}^{m}\right)-\frac{1}{\beta}\partial_{a}\left(\epsilon^{acb}e_{bp}\right)-\frac{1}{\kappa}\epsilon^{i}{}_{kp}E_{i}^{c}J^{k}\right]$
@@
$\mathcal{G}^{\text{QED}}=\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\mathcal{J}^{0}\,.$ (143)
@@
4.3 Minimal vs. non-minimal coupling
So far, our analysis has been based on the assumption of minimal coupling between gravity and fermionic fields. However, some authors have pointed out that this procedure may not be adequate in curved spacetime Bojowald:2010qpa ; PhysRevD.73.084016 ; Alexandrov:2008iy . In particular, they found that employing minimal coupling can result in a system that lacks parity invariance. As noted in Das:2008r , this issue may stem from sign inconsistencies or the choice of representation for the gamma matrices. Although, in our work we did not encounter this problem, nevertheless, for the sake of completeness, we now examine how our earlier results are affected when non-minimal coupling is assumed, focusing specifically on the case of Weyl fermions. Generalization to Dirac spinors is straightforward.
To motivate the need for this modification, we begin by highlighting a potentially problematic term in the minimally coupled gravity-fermion action:
$S_{\text{G+F}}=\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}F_{ab}^{KL}(\omega)+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\Bigr{\}}\,.$ (146)
As pointed out before, the presence of fermions modifies the compatibility condition. The difference between the full covariant derivative and the torsion-free one can be expressed as
$(\mathcal{D}_{a}-\tilde{\mathcal{D}}_{a})X_{I}=C_{aI}\,^{J}X_{J}$ (147)
where $C_{a}^{IJ}$ is the contorsion tensor. The corresponding curvature then becomes
$F_{ab}^{IJ}=\tilde{F}_{ab}^{IJ}+2\partial_{[a}C_{b]}^{IJ}+[C_{a},C_{b}]^{IJ}$ (148)
Substituting this expression into the action $S_{\text{G+F}}$, we obtain
$\begin{split}S_{\text{G+F}}=&\tilde{S}_{\text{G}}+\tilde{S}_{\text{F}}+S_{\text{int}}\\ =&\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}\tilde{F}_{ab}^{KL}(\tilde{\omega})+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}\tilde{\mathfrak{D}}_{a}\Psi-e_{I}^{a}\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{I}\Psi\Bigr{\}}\\ &+\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}[C_{a},C_{b}]^{KL}-\frac{1}{4}\int d^{4}x|\text{det}\,e|\epsilon^{I}\,_{JKL}e^{a}_{I}C_{a}^{JK}J^{L}\,.\end{split}$ (149)
Here, $\tilde{S}_{\text{G}}$ and $\tilde{S}_{\text{F}}$ are respectively the gravitational and fermionic actions constructed using the torsion-free spin connection 1-form $\tilde{\omega}_{a}^{IJ}$ and $S_{\text{int}}$ is the minimal coupling interaction term between gravity and fermions. After performing the contractions in the interaction terms, the full action becomes:
$S_{\text{G+F}}=\tilde{S}_{\text{G}}+\tilde{S}_{\text{F}}-\frac{3\kappa}{16}\frac{\beta^{2}}{1+\beta^{2}}\int d^{4}x|\text{det}\,e|J_{I}J^{I}.$ (150)
This coupling term explicitly depends on the Barbero-Immirzi parameter $\beta$, which some authors believe would be problematic PhysRevD.73.084016 . It is thus suggested that non-minimal couplings could provide a more consistent framework since setting the coupling constant $\alpha=-\frac{1}{\beta}$ yields a $\beta$-independent interaction term. Hence, we present the idea behind this suggestion and the corresponding modifications to the constraints and compatibility conditions when a non-minimal interaction is introduced.
The action incorporating a non-minimal coupling of gravity to Weyl fermions can be written as
$\begin{split}S_{\text{G+F}}=&\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}F_{ab}^{KL}(\omega)\\ &+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{I}\Psi\Bigr{\}}\,,\end{split}$ (151)
where $\alpha\in\mathbb{C}$ parametrizes the non-minimal coupling. Following the same procedure as discussed in detail in previous sections, the constraints become
$\mathcal{G}_{i}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}+\frac{i}{2}(\alpha-\bar{\alpha})\sqrt{\text{det}\,h}\mathcal{J}_{i},$ (152)
$\mathcal{S}_{i}^{\text{G+F}}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}\,J_{i}-\frac{1}{4}(-\alpha+\bar{\alpha})\sqrt{\text{det}\,h}J_{i}$
$-\frac{i\beta}{2}(\alpha-\bar{\alpha})\sqrt{\text{det}\,h}\mathcal{J}_{i},$ (153)
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{i}\Psi\right),$ (154)
$\mathcal{H}_{a}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\left(E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\right)$
$-\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi+\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{0}\Psi\right).$ (155)
In particular case of a real coupling $\alpha=\bar{\alpha}$, the constraints simplify and the new compatibility condition becomes
$\begin{split}\mathcal{D}_{b}(|e|e_{K}^{[a}e_{L}^{b]})=&\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}|e|\Bigl{[}\left(1-\frac{\alpha}{\beta}\right)\epsilon_{KL}{}^{MN}e_{M}^{a}J_{N}-\left(\frac{1}{\beta}+\alpha\right)\left(e_{K}^{a}J_{L}-e_{L}^{a}J_{K}\right)\Bigr{]}.\end{split}$ (156)
Consequently, the spin connection is modified as follow
$\Gamma_{a}^{i}=\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\Bigl{[}\left(1-\frac{\alpha}{\beta}\right)e_{a}^{i}J^{0}-\left(\frac{1}{\beta}+\alpha\right)\epsilon^{i}\,_{jk}e_{a}^{j}J^{k}\Bigr{]}.$ (157)
Finally, imposing the new compatibility condition on the Gauss constraint leads to
$\mathcal{G}_{i}^{\text{G+F}}=\frac{1}{\kappa}\epsilon_{ij}\,^{k}K_{a}^{j}E_{k}^{a}+\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\left(1-\frac{\alpha}{\beta}\right)\sqrt{\text{det}\,h}J_{i}$ (158)

### v2

@@ 3 Fermionic Field
4 Canonical Electrodynamics
@@ 4.1 The full Hamiltonian
In this section, we examine a system in which photons —denoted by a $\gamma$ with no indices— are coupled to a fermionic field, and both fields are minimally coupled to gravity. Unlike the classical formulation of canonical electrodynamics on a fixed flat spacetime, here the background is allowed to be dynamical. There have been several attempts to couple the gravitational field to electromagnetism Das (2008); Bojowald (2010). However, to the best of our knowledge, no work has considered a framework in which gravity, fermions, and photons are all simultaneously coupled. In addition, in Das (2008) a specific choice for the shift vector is adopted, which differs from the standard decomposition. As a consequence, certain time-derivative terms do not cancel in the intermediate steps of the derivation, leading to an extra contribution in the electromagnetic sector of the Hamiltonian.
To maintain clarity and avoid an overload of calculations at once, we begin by analyzing the CED action in curved spacetime using modified Ashtekar-Barbero variable. After deriving the CED contributions to the constraints, we will then proceed to minimally couple the full system to the gravitational field.
The CED action in curved spacetime takes the form:
$S_{\text{CED}}=$ $S_{F}+S_{\gamma}$ (71)
@@
$\times(\nabla_{c}A_{d}-\nabla_{d}A_{c})$
@@
$\times\left(\dot{A}_{d}-\nabla_{d}(t^{c}A_{c})-N^{c}F_{cd}\right)\,.$
@@
Let us now perform the same analysis for the full CED action $S_{\text{CED}}$ in (71). Beginning with the contributions involving the time vector $t^{a}$, we identify the following terms
@@
We observe that the first integral corresponds to the symplectic term, while the second integral contains two Gauss constraints, each multiplied by a Lagrange multiplier, respectively $\omega_{t}^{ij}$ and $A_{t}$. One of these constraints depends solely on internal indices and represents the fermionic contribution to the vacuum Gauss constraint. The other involves only spacetime indices —specifically, the spatial ones— and reduces to Gauss’ Law in the absence of the fermionic field. Hence, the full CED contribution to the Gauss constraint is
$\Lambda\mathcal{G}^{\text{CED}}=\omega_{t}{}^{ij}\,\frac{i}{2}\sqrt{\text{det}\,h}\,\Psi^{\dagger}\sigma_{ij}\Psi+A_{t}\left(\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\Psi^{\dagger}\Psi\right)\,.$ (82)
@@
Next, considering the terms involving the lapse function $N$, we obtain the CED Hamiltonian constraint as
$\begin{split}\mathcal{H}^{\text{CED}}&=\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{i}\Psi\right)\\ &+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,.\end{split}$ (83)
Finally, the terms multiplied by the shift vector $N^{a}$ yield the CED diffeomorphism constraint
$\mathcal{H}_{a}^{\text{CED}}=F_{ab}\pi^{b}+\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (84)
@@
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{CED}}=$ $\left(\beta\omega_{t}{}^{0i}-\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{t}{}^{jk}\right)\Bigl[\frac{1}{\kappa\beta}\left(\partial_{a}E_{i}^{a}+\epsilon_{ip}{}^{q}A_{a}^{p}E_{q}^{a}\right)+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\Bigr]$
@@
[…]
@@
[…]
@@
${+}\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+qE_{i}^{a}A_{a}\mathcal{J}^{i}+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,$, (92)
@@
[…]
which is consistent with the well-known CED EOM for fermions. On the other hand, the EOM for the photon field can be obtained by extremizing the full action $S_{\text{Full}}$ with respect to the electromagnetic 4-vector potential $A_{a}$. This yields a background-independent form of Maxwell’s equations. More concretely, the photonic Euler-Lagrange equation reads
$\frac{\delta\mathcal{L}^{\text{CED}}}{\delta A_{b}}-\nabla_{c}\frac{\delta\mathcal{L}^{\text{CED}}}{\delta\left(\nabla_{c}A_{b}\right)}=0\,$, (95)
@@
which is consistent with the CED theory.
@@
$(t^{a}-N^{a})$ $\Bigl[i\gamma^{0}D_{a}^{(A)}\xi-\frac{\beta}{2}K_{a}^{i}\gamma_{5}\gamma_{i}\xi-qA_{a}\gamma^{0}\xi\Bigr]-N\Bigl[ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi{-}\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi$
@@
Finally, once again we impose the compatibility condition and observe that the $\mathcal{S}_{i}$ constraint is proportional to the Gauss constraint arising from the fermionic modification of the vacuum. As a result, $\mathcal{S}_{i}$ can be discarded. Therefore, the remaining (first class) constraints are the Gauss constraint derived in the previous section and the one associated with the CED sector, together with the Hamiltonian and the diffeomorphism constraints. Moreover, since we are working with Dirac spinors —and more generally within the framework of CED— the considerations regarding chirality remain applicable. Consequently, the final full constraints, expressed in terms of the modified Ashtekar-Barbero connection and the half-density fermionic field, are
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{CED}}=$ $\Lambda^{i}\Bigl[\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{i}{2}\frac{\beta^{2}}{1+\beta^{2}}\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi\Bigr]+A_{t}\left(\nabla_{a}\pi^{a}+iq\pi\xi\right)\,$, (101)
[…]
@@
$-\frac{1}{8}({3}+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}({-3}+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-im\pi\gamma^{0}\xi$
@@ 4.2 Parity transformation
In this final subsection, we analyze the behavior of the full system under parity transformation. A similar study was conducted in Das (2008) for a system where gravity is coupled solely to a fermionic field. Here, we extend the analysis to the full system, where the gravitational field is minimally coupled to both fermionic and photonic fields, providing clearer, more detailed, and more readable final results.
@@
$+N\left[\frac{1}{\beta\sqrt{\text{det}\,h}}\left(E_{p}^{c}E_{m}^{a}A_{a}^{m}-E_{m}^{c}E_{p}^{a}A_{a}^{m}\right)-\frac{1}{\beta}\partial_{a}\left(\epsilon^{acb}e_{bp}\right){+}\frac{1}{\kappa}\epsilon^{i}{}_{kp}E_{i}^{c}J^{k}\right]$
@@
$\mathcal{G}^{\text{CED}}=\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\mathcal{J}^{0}\,.$ (134)